20 May 2003
Attendees: J. Acree, R. Fisher, J. Ford, R. Lacasse, C. Niday, W. Sizemore, D. Wirt
Jeff requested feedback on ways to
improve the IPG progress report for the
Discussion prompted the suggestion to educate county officials and local realtors about the WV State Zoning Code 37A-1 in an effort to maintain favorable public relations concerning any future development within a ten mile radius of Green Bank.
Rich especially wanted to thank all the people who helped with this effort. In a nutshell, they found that the cabinets were poorly shielded. In summary, a portion of the contributions follow: Screw spacing was changed to every 3 inches because it was determined that the 6-inch screw spacing was inadequate. Randy and Rich selected new filters; Nathan and Tracy performed chamber testing; Jeff Cromer designed the box. This effort seems to have paid off; L-band testing appears to be good and plans are set to test at 300 MHz this week.
Chuck Niday asked about the penetrations caused by the lockouts installed by Tim Weadon. Rich stated this installation was done properly; they used a long metal tube serving as a waveguide to eliminate RFI from passing through the installation.
Wesley stated that there is no comprehensive database on satellite frequencies. Most information we have is from amateur operator web sites. The observer or operator needs to assist us by providing as much information as possible about the RFI characteristics, such as exact sky frequency, variability, and location (i.e. moving or stationary). If the signal appears to be unintentional, then it is probably an on-site source. If it is intentional, then frequency allocation tables can be used to indicate possible user. Wesley stated that he has several contacts he can call on as well as performing web searches. There are several sites on the web that provide satellite data; however, most of the sites are specific to particular satellite constellations (i.e. Amateur, Weather, Space Shuttle/ISS, etc.)
John Ford would like for us to build a comprehensive database that the astronomers can use by plugging in a proposed frequency and generating all known sources. It is a shame that we are not providing information on sources that we are aware of. Even if it is unidentified, but exists, we can let the astronomers know about it. Ideally, we would have a combination of a searchable database and legacy data for use as resources.
Wes reported that the monitoring station is out of commission until the broadband low noise amplifier is back from Miteq, which is being repaired due to lightning damage. The spare was being used until it was also zapped this week. John approved the money for the pending repair. Although surge protection is in place, Jeff and Wes will review the monitoring station design to ensure that a problem does not exist before installing the repaired amplifier. Jeff has successfully used similar set ups in the past, and does not initially see why we are experiencing such frequent lightning damage.
John reported on the 2003 budget outlook. After initially being told to keep spending at about the same level as last year, he received his budget on Friday which reflects a cut from last year’s budget. This cut is due to money being diverted to support the GBT track work. At this point, being into the second quarter, he is consequently over-budget. His spending rules to compensate for this is: No new infrastructure; no outsourced calibration; no new projects or possible suspension of some projects; critical repairs to our operation are acceptable as well as critical RFI control. Basically we will be operating on an “emergency only” spending level.
The outlook for 2004 is much the same since work on the track will be ongoing into next year and the budget from NSF likely will not increase. The only positive note about FY04 is that we know about it now and can plan accordingly.
Wes reported that the integrity of the shielded rooms is in tact. Nathan Sharp led the effort being involved with the entire project start to finish. The shielding around the two large windows and exterior door in the OVLBI control room was finished on Friday, May 16. The only outstanding items remaining on the scope of work is to install the window moisture barrier and the flectron material, but these items should not affect the integrity of the shielding.
This item will be reported on next meeting.